The point P is on the unit circle. Find P(x,y) from the given information. The x coordinate of P is 4/5 and P is in Quadrant I.
It's been a while since I had to remember to do this. Can someone please help remind me what they are asking here. Thank you

Answer by stanbon(75887)

The point P is on the unit circle. Find P(x,y) from the given information. The x coordinate of P is 4/5 and P is in Quadrant I.
-------
Since the point is on the circumference of a unit circle,
x^2 + y^2 = 1
---------------------------
Your point is ((4/5),y) and y is positive
(4/5)^2 + y^2 = 1
y^2 = (25/25)-(16/25)
y^2 = 9/25
y = 3/5
-------------------
Cheers,
Stan H.
===================